# Lumiere: A Space-Time Diffusion Model for Video Generation

### 23 Jan 2024 v1

@@ 2 Related work
Recently, there have been substantial efforts in training large-scale T2V models on large scale datasets with autoregressive Transformers (e.g., (Villegas et al., 2023; Wu et al., 2022; Hong et al., 2022; Kondratyuk et al., 2023)) or Diffusion Models (e.g., (Ho et al., 2022a, b)). A prominent approach for T2V generation is to “inflate” a pre-trained T2I model by inserting temporal layers to its architecture, and fine-tuning only those, or optionally the whole model, on video data (Singer et al., 2022; Blattmann et al., 2023b; Girdhar et al., 2023; Ge et al., 2023; Yuan et al., 2024). PYoCo (Ge et al., 2023) carefully design video noise prior and obtain better performance for fine-tuning a T2I model for video generation. VideoLDM (Blattmann et al., 2023b) and AnimateDiff (Guo et al., 2023) inflate StableDiffusion (Rombach et al., 2022) and train only the newly-added temporal layers, showing they can be combined with the weights of personalized T2I models. Interestingly, the ubiquitous convention of existing inflation schemes is to maintain a fixed temporal resolution across the network, which limits their ability to process full-length clips. In this work, we design a new inflation scheme which includes learning to downsample the video in both space and time, and performing the majority of computation in the compressed space-time feature space of the network. We extend an Imagen T2I (Saharia et al., 2022b), however our architectural contributions could be used for latent diffusion as well, and are orthogonal to possible improvements to the diffusion noise scheduler (Ge et al., 2023) or to the video data curation (Blattmann et al., 2023a).
@@ 3.1 Space-Time U-Net (STUnet)
Similarly to (Blattmann et al., 2023b; Guo et al., 2023), we train the newly added parameters, and keep the weights of the pre-trained T2I fixed. Notably, the common inflation approach ensures that at initialization, the T2V model is equivalent to the pre-trained T2I model, i.e., generates videos as a collection of independent image samples. However, in our case, it is impossible to satisfy this property due to the temporal down- and up-sampling modules. We empirically found that initializing these modules such that they perform nearest-neighbor down- and up- sampling operations results with a good starting point (in terms of the loss function).
@@ 3.2 Multidiffusion for Spatial-Super Resolution
The solution to this problem is given by linearly combining the predictions over overlapping windows.
@@ 4.1 Stylized Generation
Figure 6 presents sample results for various styles from (Sohn et al., 2023). While more realistic styles such as “watercolor painting” result in realistic motion, other, less realistic spatial priors derived from vector art styles, result in corresponding unique non-realistic motion. For example, the “line drawing” style results in animations that resemble pencil strokes “drawing” the described scene, while the “cartoon” style result in content that gradually “pops out” and constructs the scene (see SM for full videos).
@@ 5 Evaluation and Comparisons
We train our T2V model on a dataset containing 30M videos along with their text caption. The videos are 80 frames long at 16 fps (5 seconds). The base model is trained at $128\times 128$ and the SSR outputs $1024\times 1024$ frames. We evaluate our model on a collection of 113 text prompts describing diverse objects and scenes. The prompt list consists of 18 prompts assembled by us and 95 prompts used by prior works (Singer et al., 2022; Ho et al., 2022a; Blattmann et al., 2023b) (see App. B). Additionally, we employ a zero-shot evaluation protocol on the UCF101 dataset (Soomro et al., 2012), as detailed in Sec. 5.2.
@@
Figure 10: User study results. Our method was preferred by users in both text-to-video, and image-to-video generation. See Sec. 5.2.
@@
We compare our method to prominent T2V diffusion models: (i) ImagenVideo (Ho et al., 2022a), that operates in pixel-space and consists of a cascade of 7 models (a base model, 3 TSR models, and 3 SSR models); (ii) AnimateDiff (Guo et al., 2023), (iii) StableVideoDiffusion (SVD) (Blattmann et al., 2023a), and (iv) ZeroScope (Wang et al., 2023a) that inflate Stable Diffusion (Rombach et al., 2022) and train on video data; note that AnimateDiff and ZeroScope output only 16, and 36 frames respectively. SVD released only their image-to-video model, which outputs 25 frames and is not conditioned on text. Additionally, we compare to (v) Pika (Pika labs, 2023) and (vi) Gen-2 (RunwayML, 2023) commercial T2V models that have available API. Furthermore, we qualitatively compare to additional T2V models that are closed-source in Sec. 5.2.
@@ 5.2 Quantitative Evaluation
We further conduct a user study for comparing our image-to-video model (see Sec. 4.2) against Pika (Pika labs, 2023), StableVideoDiffusion (SVD) (Blattmann et al., 2023a), and Gen2(RunwayML, 2023). Note that SVD image-to-video model is not conditioned on text, we thus focus our survey on the video quality. As seen in Fig. 10, our method was preferred by users compared to the baselines. For a detailed description of the full evaluation protocol, please refer to Appendix B.
@@ 7 Societal Impact
Our primary goal in this work is to enable novice users to generate visual content in an creative and flexible way. However, there is a risk of misuse for creating fake or harmful content with our technology, and we believe that it is crucial to develop and apply tools for detecting biases and malicious use cases in order to ensure a safe and fair use.
@@ Acknowledgments
We would like to thank Miki Rubinstein, Amit Raj, Ronny Votel, Orly Liba, Bryan Seybold, David Ross, Guanghui Liu, Dan Goldman, Hartwig Adam, Xuhui Jia, Xiuye Gu, Rachel Hornung, Oran Lang and Jess Gallegos for their collaboration, helpful discussions, feedback and support. We thank owners of images and videos used in our experiments for sharing their valuable assets (attributions can be found in our webpage).
@@ Appendix A Qualitative Comparison
Figure 11: Qualitative Comparison. We qualitatively compare our model to prominent T2V diffusion models, Gen-2 (RunwayML, 2023), Pika (Pika labs, 2023), ImagenVideo (Ho et al., 2022a), AnimateDiff (Guo et al., 2023) and ZeroScope (Wang et al., 2023a).
Appendix B Evaluation Protocol
B.1 User study
@@
We used a set of 113 text prompts containing a variety of objects and actions for the text-to-video and image-to-video surveys. $95$ prompts were taken from various recent text-to-video generation methods (Singer et al., 2022; Ho et al., 2022a; Blattmann et al., 2023b). The rest are prompts we created which describing complex scenes and actions, see Sec. B.3 for the full list of prompts. For each baseline method, we generated results for all prompts through their official APIs. To ensure a fair comparison, when generating results using our method we’ve fixed a random seed and generated all results using the same random seed without any curation. In addition, following (Girdhar et al., 2023) we align the spatial resolution and temporal duration of our videos to each one of the baselines. Spatially, we apply a central crop followed by resizing to 512$\times$512, and temporally, by trimming the start and the end such that the duration of both videos matches the shortest one. Each participant is shown 10 side-by-side comparisons between our method and a baseline (randomly ordered). We then asked the participant ”Which video matches the following text?” and ”Which video is better? Has more motion and better quality”. We enable the user to answer only after watching the full video, waiting at least 3 seconds, and pressing a button reading “Ready to answer” (Fig. 12). To ensure credible responses, we include several vigilance tests during the study - a real video vs a complete random noise, and a static image vs a real video.
@@
B.2 Zero-shot evaluation on UCF101
@@
B.3 List of prompts used for the user study evaluation
@@
    ”A storm trooper riding a bike in sunset street, 4k, high resolution.”
@@
    ”A storm trooper vacuuming the beach.”
@@
    ”A swarm of bees flying around their hive.”
@@
    ”A teddy bear is playing the electric guitar, high definition, 4k.”
@@
    ”A teddy bear running in New York City.”
@@
    ”A video of the Earth rotating in space.”
@@
    ”Aerial view of a hiker man standing on a mountain peak.”
@@
    ”Aerial view of a snow-covered mountain.”
@@
    ”An animated painting of fluffy white clouds moving in sky.”
@@
    ”An astronaut riding a horse, high definition, 4k.”
@@
    ”An astronaut riding a horse in sunset, 4k, high resolution.”
@@
    ”Beer pouring into glass, low angle video shot.”
@@
    ”Bird-eye view of a highway in Los Angeles.”
@@
    ”Cartoon character dancing.”
@@
    ”Cherry blossoms swing in front of ocean view, 4k, high resolution.”
@@
    ”Chocolate syrup pouring on a vanilla ice cream”
@@
    ”Close up of grapes on a rotating table. High definition.”
@@
    ”Teddy bear surfer rides the wave in the tropics.”
@@
    ”Drone flythrough interior of sagrada familia cathedral.”
@@
    ”Drone flythrough of a tropical jungle covered in snow.”
@@
    ”Flying through fantasy landscapes, 4k, high resolution.”
@@
    ”Horror house living room interior overview design, Moebius, Greg Rutkowski, Zabrocki, Karlkka, Jayison Devadas, Phuoc Quan, trending on Artstation, 8K, ultra wide angle, pincushion lens effect.”
@@
    ”Hulk wearing virtual reality goggles, 4k, high resolution.”
@@
    ”Ironman flying over a burning city, very detailed surroundings, cities are blazing, shiny iron man suit, realistic, 4k ultra high definition.”
@@
    ”Jack russell terrier dog snowboarding. GoPro shot.”
@@
    ”Low angle of pouring beer into a glass cup.”
@@
    ”Melting pistachio ice cream dripping down the cone.”
@@
    ”Milk dripping into a cup of coffee, high definition, 4k.”
@@
    ”Pouring latte art into a silver cup with a golden spoon next to it.
@@
    ”Time lapse at a fantasy landscape, 4k, high resolution.”
@@
    ”Tiny plant sprout coming out of the ground.”
@@
    ”Toy Poodle dog rides a penny board outdoors”
@@
    ”Traveler walking alone in the misty forest at sunset.”
@@
    ”Turtle swimming in ocean.”
@@
    ”Two pandas sitting at a table playing cards, 4k, high resolution.”
@@
    ”Two raccoons reading books in NYC Times Square.”
@@
    ”Waves crashing against a lone lighthouse, ominous lighting.”
  110. 110.
    ”Woman in white dress waving on top of a mountain.”
  111. 111.
    ”Wood on fire.”
  112. 112.
    ”Wooden figurine surfing on a surfboard in space.”
  113. 113.

### 05 Feb 2024 v2

@@ 2 Related work
Recently, there have been substantial efforts in training large-scale T2V models on large scale datasets with autoregressive Transformers (e.g., (Villegas et al., 2023; Wu et al., 2022; Hong et al., 2022; Kondratyuk et al., 2023)) or Diffusion Models (e.g., (Ho et al., 2022a, b; Gupta et al., 2023)). A prominent approach for T2V generation is to “inflate” a pre-trained T2I model by inserting temporal layers to its architecture, and fine-tuning only those, or optionally the whole model, on video data (Singer et al., 2022; Blattmann et al., 2023b; Girdhar et al., 2023; Ge et al., 2023; Yuan et al., 2024). PYoCo (Ge et al., 2023) carefully design video noise prior and obtain better performance for fine-tuning a T2I model for video generation. VideoLDM (Blattmann et al., 2023b) and AnimateDiff (Guo et al., 2023) inflate StableDiffusion (Rombach et al., 2022) and train only the newly-added temporal layers, showing they can be combined with the weights of personalized T2I models. Interestingly, the ubiquitous convention of existing inflation schemes is to maintain a fixed temporal resolution across the network, which limits their ability to process full-length clips. In this work, we design a new inflation scheme which includes learning to downsample the video in both space and time, and performing the majority of computation in the compressed space-time feature space of the network. We extend an Imagen T2I model (Saharia et al., 2022b), however our architectural contributions could be used for latent diffusion as well, and are orthogonal to possible improvements to the diffusion noise scheduler (Ge et al., 2023) or to the video data curation (Blattmann et al., 2023a).
@@ 3.1 Space-Time U-Net (STUnet)
Similarly to (Blattmann et al., 2023b; Guo et al., 2023), we train the newly added parameters, and keep the weights of the pre-trained T2I fixed. Notably, the common inflation approach ensures that at initialization, the T2V model is equivalent to the pre-trained T2I model, i.e., generates videos as a collection of independent image samples. However, in our case, it is impossible to satisfy this property due to the temporal down- and up-sampling modules. We empirically found that initializing these modules such that they perform nearest-neighbor down- and up- sampling operations results with a good starting point (see App. B).
@@ 3.2 Multidiffusion for Spatial-Super Resolution
The solution to this problem is given by linearly combining the predictions over overlapping windows. See App. C.
@@ 4.1 Stylized Generation
Figure 6 presents sample results for various styles from (Sohn et al., 2023). While more realistic styles such as “watercolor painting” result in realistic motion, other, less realistic spatial priors derived from vector art styles, result in corresponding unique non-realistic motion. For example, the “line drawing” style results in animations that resemble pencil strokes “drawing” the described scene, while the “cartoon” style results in content that gradually “pops out” and constructs the scene (see SM for full videos).
@@ 5 Evaluation and Comparisons
We train our T2V model on a dataset containing 30M videos along with their text caption. The videos are 80 frames long at 16 fps (5 seconds). The base model is trained at $128\times 128$ and the SSR outputs $1024\times 1024$ frames. We evaluate our model on a collection of 109 text prompts describing diverse objects and scenes. The prompt list consists of 91 prompts used by prior works (Singer et al., 2022; Ho et al., 2022a; Blattmann et al., 2023b) and the rest were created by us (see App. D). Additionally, we employ a zero-shot evaluation protocol on the UCF101 dataset (Soomro et al., 2012), as detailed in Sec. 5.2.
@@
Figure 10: User study. We compare our method to each of the baselines. For each baseline, we report the percentage of user votes in our favor (blue) and in favor of the baseline (orange). Our method was preferred by users in both text-to-video, and image-to-video generation. See Sec. 5.2.
@@
We compare our method to prominent T2V diffusion models: (i) ImagenVideo (Ho et al., 2022a), that operates in pixel-space and consists of a cascade of 7 models (a base model, 3 TSR models, and 3 SSR models); (ii) AnimateDiff (Guo et al., 2023), (iii) StableVideoDiffusion (SVD) (Blattmann et al., 2023a), and (iv) ZeroScope (Wang et al., 2023a) that inflate Stable Diffusion (Rombach et al., 2022) and train on video data; note that AnimateDiff and ZeroScope output only 16, and 36 frames respectively. SVD released only their image-to-video model, which outputs 25 frames and is not conditioned on text. Additionally, we compare to (v) Pika (Pika labs, 2023) and (vi) Gen-2 (RunwayML, 2023) commercial T2V models that have available API. Furthermore, we quantitatively compare to additional T2V models that are closed-source in Sec. 5.2.
@@ 5.2 Quantitative Evaluation
We further conduct a user study for comparing our image-to-video model (see Sec. 4.2) against Pika (Pika labs, 2023), StableVideoDiffusion (SVD) (Blattmann et al., 2023a), and Gen2(RunwayML, 2023). Note that SVD image-to-video model is not conditioned on text, we thus focus our survey on the video quality. As seen in Fig. 10, our method was preferred by users compared to the baselines. For a detailed description of the full evaluation protocol, please refer to Appendix D.
@@ 7 Societal Impact
Our primary goal in this work is to enable novice users to generate visual content in a creative and flexible way. However, there is a risk of misuse for creating fake or harmful content with our technology, and we believe that it is crucial to develop and apply tools for detecting biases and malicious use cases in order to ensure a safe and fair use.
@@ Acknowledgments
We would like to thank Ronny Votel, Orly Liba, Hamid Mohammadi, April Lehman, Bryan Seybold, David Ross, Dan Goldman, Hartwig Adam, Xuhui Jia, Xiuye Gu, Mehek Sharma, Rachel Hornung, Oran Lang, Jess Gallegos, William T. Freeman and David Salesin for their collaboration, helpful discussions, feedback and support. We thank owners of images and videos used in our experiments for sharing their valuable assets (attributions can be found in our webpage).
@@ Appendix A Qualitative Comparison
Figure 11: Qualitative Comparison. We provide samples from our user study, comparing our model to prominent T2V diffusion models, Gen-2 (RunwayML, 2023), Pika (Pika labs, 2023), ImagenVideo (Ho et al., 2022a), AnimateDiff (Guo et al., 2023) and ZeroScope (Wang et al., 2023a).
Appendix B Initialization
As illustrated in Fig. 4(b)-(c), each inflation block has a residual structure, in which the temporal components are followed by a linear projection, and are added to the output of the pre-trained T2I spatial layer. Similarly to previous works (e.g., (Guo et al., 2023)), we initialize the linear projection to a zero mapping, such that at initialization, the inflation block is equivalent to the pre-trained T2I block. However, as mentioned in Sec. 3.1, due to the temporal down- and up- sampling modules, the overall inflated network cannot be equivalent at initialization to the pre-trained T2I model due to the temporal compression. We empirically found that initializing the temporal down- and up-sampling modules such that they perform nearest-neighbor down- and up- sampling operations (i.e., temporal striding or frame duplication) is better than a standard initialization (He et al., 2015). In more detail, the temporal downsampling is implemented as a 1D temporal convolution (identity initialized) followed by a stride, and the temporal upsampling is implemented as frame duplication followed by a 1D temporal convolution (identity initialized). Note that this scheme ensures that at initialization, every $N^{\text{th}}$ frame is identical to the output of the pre-trained T2I model, where $N$ is the total temporal downsampling factor in the network (see Fig. 13). We ablate our choice of temporal down- and up- sampling initialization on the UCF-101 (Soomro et al., 2012) dataset in Fig. 12.
Figure 12: Initialization Ablation. We ablate the initialization of our temporal down- and up-sampling modules by training on UCF-101 (Soomro et al., 2012) and comparing the loss function. See App. B for details.
Figure 13: Random Samples at Initialization. We visualize several frames generated at initialization with different initialization schemes for the temporal down- and up- sampling modules, to demonstrate their effect. The standard random initialization (first row) results with non-meaningful samples, because the temporal down- and up-sampling modules mix unrelated feature maps. The identity initialization (second row) allows us to better benefit from the prior of the pre-trained text-to-image. See App. B for details.
Appendix C SSR with MultiDiffusion
As discussed in Sec. 3.2, we apply MultiDiffusion (Bar-Tal et al., 2023) along the temporal axis in order to feed segments of the video to the SSR network. Specifically, the temporal windows have an overlap of 2 frames, and at each denoising step we average the predictions of overlapping pixels. MultiDiffusion allows us to avoid temporal boundary artifacts between segments of the video (see Fig. 14).
Figure 14: MultiDiffusion Ablation. We ablate the use of MultiDiffusion (Bar-Tal et al., 2023) for the SSR network. We show sample high-resolution SSR samples in the first row. Below, we show xt-slice of our result and of of the SSR network applied without MultiDiffusion. As can be seen, without MultiDiffusion, there are temporal boundary artifcats. We also demonstarte this property by computing Frame Diff. Specifically, we compute the squared difference between adjacent frames. Without MultiDiffusion, there is a high peak of error between every non-overlapping segment. See App. C.
Appendix D Evaluation Protocol
D.1 User study
@@
We used a set of 109 text prompts containing a variety of objects and actions for the text-to-video and image-to-video surveys. $91$ prompts were taken from various recent text-to-video generation methods (Singer et al., 2022; Ho et al., 2022a; Blattmann et al., 2023b). The rest are prompts we created which describe complex scenes and actions, see Sec. D.3 for the full list of prompts. For each baseline method, we generated results for all prompts through their official APIs. To ensure a fair comparison, when generating results using our method we’ve fixed a random seed and generated all results using the same random seed without any curation. In addition, following (Girdhar et al., 2023) we align the spatial resolution and temporal duration of our videos to each one of the baselines. Spatially, we apply a central crop followed by resizing to 512$\times$512, and temporally, by trimming the start and the end such that the duration of both videos matches the shortest one. Each participant is shown 10 side-by-side comparisons between our method and a baseline (randomly ordered). We then asked the participant ”Which video matches the following text?” and ”Which video is better? Has more motion and better quality”. We enable the user to answer only after watching the full video, waiting at least 3 seconds, and pressing a button reading “Ready to answer” (Fig. 15). To ensure credible responses, we include several vigilance tests during the study - a real video vs a complete random noise, and a static image vs a real video.
@@
D.2 Zero-shot evaluation on UCF101
@@
D.3 List of prompts used for the user study evaluation
@@
    ”A swarm of bees flying around their hive.”
@@
    ”A teddy bear is playing the electric guitar, high definition, 4k.”
@@
    ”A teddy bear running in New York City.”
@@
    ”A teddy bear skating in Times Square.”
@@
    ”A video of the Earth rotating in space.”
@@
    ”Aerial view of a hiker man standing on a mountain peak.”
@@
    ”Aerial view of a snow-covered mountain.”
@@
    ”An animated painting of fluffy white clouds moving in sky.”
@@
    ”An astronaut riding a horse, high definition, 4k.”
@@
    ”An astronaut riding a horse in sunset, 4k, high resolution.”
@@
    ”Beer pouring into glass, low angle video shot.”
@@
    ”Cartoon character dancing.”
@@
    ”Cherry blossoms swing in front of ocean view, 4k, high resolution.”
@@
    ”Chocolate syrup pouring on a vanilla ice cream”
@@
    ”Close up of grapes on a rotating table. High definition.”
@@
    ”Teddy bear surfer rides the wave in the tropics.”
@@
    ”Drone flythrough interior of sagrada familia cathedral.”
@@
    ”Drone flythrough of a tropical jungle covered in snow.”
@@
    ”Flying through fantasy landscapes, 4k, high resolution.”
@@
    ”Horror house living room interior overview design, 8K, ultra wide angle, pincushion lens effect.”
@@
    ”Jack russell terrier dog snowboarding. GoPro shot.”
@@
    ”Low angle of pouring beer into a glass cup.”
@@
    ”Melting pistachio ice cream dripping down the cone.”
@@
    ”Milk dripping into a cup of coffee, high definition, 4k.”
@@
    ”Pouring latte art into a silver cup with a golden spoon next to it.
@@
    ”Shoveling snow.”
@@
    ”Time lapse at a fantasy landscape, 4k, high resolution.”
@@
    ”Traveler walking alone in the misty forest at sunset.”
@@
    ”Two pandas discussing an academic paper.”
@@
    ”Two pandas sitting at a table playing cards, 4k, high resolution.”
@@
    ”Two raccoons reading books in NYC Times Square.”
@@
    ”Waves crashing against a lone lighthouse, ominous lighting.”
@@
    ”Woman in white dress waving on top of a mountain.”